Three point charges are located at the following positions Q1 200 uC at x -1.00 m 02 300 uC atx 0,Q3 500 uC atx-100m What is the magnitude of the lorce on the 3 00-pC charge? 810 102 N 5.40 10 2 N O 0.135 N 0.189 N 0.158 N
Answer #1

Electrostatic force is given by:

F = k*q1*q2/r^2

Force will be attractive if both charge has opposite sign and Force will be repulsive if both charge has same signs,

Now net force on Charge Q2 will be

F12 (force due to charge Q1 on charge Q2) will be repulsive force and towards the -ve x-axis

F23 (force due to charge Q3 on charge Q2) will be attractive force and towards the -ve x-axis

So, Net force will be

F2 = -F12 - F23

Magnitude of net force will be

|F2| = |F12 + F23|

So,

|F2| = k*Q1*Q2/r12^2 + k*Q2*Q3/r23^2

r12 = distance between Q1 and Q2 = 1.00 - 0 = 1.00 m

r23 = distance between Q2 and Q3 = 0 - (-1.00) = 1.00 m

|F2| = 9*10^9*2*10^-6*3*10^-6/1.00^2 + 9*10^9*5*10^-6*3*10^-6/1.00^2

|F2| = 0.189 N

Correct option is D.

Please Upvote.

Comment below if you have any query.